### v1

@@ Abstract.
Audio deepfake detection (ADD) has grown increasingly important due to the rise of high-fidelity audio generative models and their potential for misuse. Given that audio large language models (ALLMs) have made significant progress in various audio processing tasks, a heuristic question arises: Can ALLMs be leveraged to solve ADD?. In this paper, we first conduct a comprehensive zero-shot evaluation of ALLMs on ADD, revealing their ineffectiveness in detecting fake audio. To enhance their performance, we propose $\mathcal{A}LLM4ADD$, an ALLM-driven framework for ADD. Specifically, we reformulate ADD task as an audio question answering problem, prompting the model with the question: “Is this audio fake or real?”. We then perform supervised fine-tuning to enable the ALLM to assess the authenticity of query audio. Extensive experiments are conducted to demonstrate that our ALLM-based method can achieve superior performance in fake audio detection, particularly in data-scarce scenarios. As a pioneering study, we anticipate that this work will inspire the research community to leverage ALLMs to develop more effective ADD systems.
@@ 1. Introduction
Recently, Audio Large Language Models (ALLMs) (Gong et al., 2024; Zhang et al., 2023b; Chu et al., 2023, 2024) have demonstrated remarkable progress across a wide range of audio processing tasks, including audio captioning (Tang et al., 2024; Shan et al., 2025; Wu et al., 2024) and speech recognition (Li et al., 2024b; Fathullah et al., 2024; Chu et al., 2023). Notable models include Qwen-Audio (Chu et al., 2023), which integrates the Whisper encoder (Radford et al., 2023) with the text Qwen LLM (Bai et al., 2023), enabling the latter to understand audio. However, their performance on ADD task remains unexplored. This raises a critical question: Can ALLMs effectively address the ADD task?
To answer this question, we present pioneering work that leverages ALLMs for ADD task. To the best of our knowledge, this is the first paper to tackle ADD using the ALLM-based approach. First, we conduct a comprehensive quantitative evaluation of ALLM’s capabilities in ADD task. Our experimental results reveal that existing ALLMs perform poorly in zero-shot fake audio detection, primarily due to the mismatch between their pretraining objectives and the fake audio detection requirements. To enhance their performance for fake audio detection, we further propose a novel framework called $\mathcal{A}LLM4ADD$. Specifically, we reformulate ADD task as an Audio Question Answering (AQA) problem, prompting the model with the question “Is this audio fake or real?” and instructing it to generate the correct answer. We then employ supervised fine-tuning (SFT) to endow the ALLMs the capability to answer “Fake” if the query audio is fake, and conversely, “Real” if it is real. Extensive experimental results demonstrate that $\mathcal{A}LLM4ADD$ achieves superior performance compared to existing conventional pipeline and end-to-end models, particularly in data-scarce scenarios. These findings collectively highlight the advantages of our ALLM-based approach for fake audio detection.
@@
  • We propose $\mathcal{A}LLM4ADD$, a novel framework which reformulates ADD task as an AQA problem, to successfully endow ALLMs with fake audio detection capability.
@@ 2.2. Audio Large Language Model
In the past year, modern Large Language Models (LLMs) have demonstrated powerful reasoning and understanding abilities (Touvron et al., 2023; Bai et al., 2023). To extend the application scope of LLMs beyond pure text tasks, many LLM-based multimodal models (Liu et al., 2023a; Bai et al., 2025; Zhang et al., 2023b; Gong et al., 2024; Zhang et al., 2023a; Sun et al., 2025) have been developed.
@@
While existing ALLMs have achieved notable success in tasks such as audio caption (Tang et al., 2024; Shan et al., 2025; Wu et al., 2024) and speech recognition (Li et al., 2024b; Fathullah et al., 2024; Chu et al., 2023), their application in detecting fake audio remains unexplored. Therefore, in this paper, we formulate audio deepfake detection as an audio question answering task, leveraging the advanced understanding capabilities of these ALLMs to address this emerging challenge.
3. Can ALLMs detect fake audio in a zero-shot manner?
@@ 4.1. Task Formulation
In order to take advantage of the Audio Large Language Model (ALLM), $\mathcal{A}LLM4ADD$ formulates the audio deepfake detection (ADD) task as a audio question answering (AQA) problem. In this framework, the input comprises two crucial components: a query audio $A$ that needs to be classified as real or fake and an instruction prompt $q$, which guides $\mathcal{A}LLM4ADD$ in its analysis of the query audio. The instruction $q$ can take on various forms (e.g., “Is this audio fake or real?”). The output of this framework corresponds to the answer text $y$. While $y$ can be any text in principle, we constrain it to two options: ”Fake” and ”Real” during training, aligning with the ground truth to the original binary classification problem.
@@ 4.2. Model Architecture
As shown in Figure 1, $\mathcal{A}LLM4ADD$ adopts an architecture similar to the Qwen-Audio series, consisting of an audio encoder and a large language model. This framework takes two inputs: a query audio and a natural language instruction. Below, we describe each sub-module in detail.
@@
Our $\mathcal{A}LLM4ADD$ adopts a LLM as its foundation component. The model is initialized with pre-trained weights from Qwen-audio-chat default, which is a 32-layer Transformer decoder model with a hidden size of 4096, encompassing a total of 7.7B parameters. The LLM is employed to process audio representations and corresponding instructions, subsequently generating responses capable of discerning authenticity. The format input to the LLM follows this format:
@@ 4.3. Supervised Fine-tuning
Additionally, we categorize our methods based on the trainability of the audio encoder: $\mathbf{\mathcal{A}LLM4ADD^{\star}}$ denotes that the encoder is frozen, while $\mathbf{\mathcal{A}LLM4ADD^{\triangle}}$ indicates that the encoder is trainable. Our motivation is to transform the feature space of the audio embeddings into one that can effectively discriminate between real and fake by training the audio encoder. This strategy aims to capture more detailed nuances of audio authenticity, thereby enhancing the model’s performance in detecting fake audio.
@@ 5.1.2. Baselines
For end-to-end models, we select five competitive methods that provide open-source code: RawNet2 (Tak et al., 2021b), AASIST (Jung et al., 2021), RawGAT-ST (Tak et al., 2021a), Rawformer (Liu et al., 2023b) and RawBMamba (Chen et al., 2024b). Brief descriptions are provided below. RawNet2 (Tak et al., 2021b) is a convolutional neural network operating directly on raw audio waveforms, utilizing residual blocks and Sinc-Layers (Ravanelli and Bengio, 2018) as band-pass filters for effective spoofed audio detection. AASIST (Jung et al., 2021) employs a heterogeneous stacking graph attention layer to model artifacts across temporal and spectral segments. RawGAT-ST (Tak et al., 2021a) utilizes spectral and temporal sub-graphs integrated with a graph pooling strategy, effectively processing complex auditory environments. Rawformer (Liu et al., 2023b) integrates convolution layer and transformer to model local and global artefacts and relationship directly on raw audio. RawBMamba (Chen et al., 2024b) proposes an end-to-end bidirectional state space model to capture both short- and long-range discriminative information.
@@ 5.1.3. Evaluation Metric
Table 3. Performance comparison of our methods with conventional pipeline and end-to-end models. (Train) and (Frozen) represent whether the self-supervised features are trainable. $\mathbf{\mathcal{A}LLM4ADD^{\star}}$ and $\mathbf{\mathcal{A}LLM4ADD^{\triangle}}$ denote the audio encoder is frozen and trainable, respectively. We train the models across: ASV@full, ASV@1/4, ASV@1/8, and ASV@1/16, and report EER (), ACC (), and AUC () on the ASVspoof2019 LA evaluation set. Best results in each column are highlighted in bold.
@@
$\mathcal{A}LLM4ADD^{\star}$ 1.30 99.26 99.88 1.97 98.87 99.65 2.46 $\mathbf{98.69}$ 99.54 3.13 98.11 $\mathbf{99.43}$
$\mathcal{A}LLM4ADD^{\triangle}$ $\mathbf{0.41}$ $\mathbf{99.39}$ $\mathbf{99.97}$ $\mathbf{1.15}$ $\mathbf{99.12}$ $\mathbf{99.71}$ $\mathbf{1.63}$ 98.55 $\mathbf{99.75}$ $\mathbf{2.45}$ $\mathbf{98.52}$ 99.39
@@ 5.2. Implementation Details
We use Qwen-audio-chat weights as our default initial weights. We optimize our model using the Adam optimizer with hyperparameters $\beta=(0.9,0.95)$ and implement a cosine learning rate scheduler. The warm-up ratio is set at 0.01 for $ASV@full$, $ASV@1/4$, and $ASV@1/8$ configurations, and 0.05 for $ASV@1/16$. We investigate two versions of our $\mathcal{A}LLM4ADD$: $\mathcal{A}LLM4ADD^{\star}$ and $\mathcal{A}LLM4ADD^{\triangle}$. For $\mathcal{A}LLM4ADD^{\star}$, the audio encoder is frozen; for $\mathcal{A}LLM4ADD^{\triangle}$, the audio encoder is trainable. The initial learning rate is determined through beam search within the range of [3e-5, 4e-5, 5e-5, 1e-4]. Additionally, we apply a weight decay of 0.1 and a gradient clipping threshold of 1.0 to maintain training stability. Training epochs varies: 5 epochs for $ASV@full$, $ASV@1/4$, and $ASV@1/8$, and increase to 10 epochs for $ASV@1/16$. To optimize GPU memory usage, we employ a gradient checkpointing technique during training. Furthermore, our training processes utilize bfloat16 precision with automatic mixed precision for efficiency. We conduct all experiments on a single Nvidia A100.
For LoRA configuration, we configure our LoRA hyperparameter as follows: LoRA rank $r$ as 64, LoRA alpha (scaling factor) as 16 and LoRA dropout as 0.05. We apply LoRA to the query, key, value, and output projection layers.
For conventional pipeline baselines, we adhere to the hyperparameter provided in (Wang and Yamagishi, 2022). For self-supervised features W2V and Hubert, we utilize checkpoints from: Wav2vec2-Base [^5] https://huggingface.co/facebook/wav2vec2-base, and Hubert-Base [^6] https://huggingface.co/facebook/hubert-base-ls960, respectively. For end-to-end baselines, we adhere to the official codebase and train the models for 100 epochs. To maintain a fair comparison across all methods, we do not employ data augmentation techniques.
@@ 5.3. Experimental Results
To demonstrate the superiority of our ALLM-based fake audio detection method, we compare it with extensive baselines detailed in Sec. 5.1.2. The models are trained on $ASV@full$, $ASV@1/4$, $ASV@1/8$, and $ASV@1/16$, and are evaluated on the ASVspoof2019 evaluation set. The experimental results are presented in Table 3. From Table 3, we can draw the following observations:
  • Our $\mathcal{A}LLM4ADD$ can achieve excellent results in ADD task. When trained on $ASV@full$, $\mathcal{A}LLM4ADD^{\triangle}$ achieves an EER of 0.41%, surpassing the best performances of end-to-end and conventional pipeline baselines, which are 0.83% and 1.94%, respectively. Furthermore, $\mathcal{A}LLM4ADD$ maintains strong performance when data is scarce. Specifically, under the $ASV@1/16$ setting, $\mathcal{A}LLM4ADD^{\triangle}$ still achieves an EER of 2.45% and an accuracy of 98.52%. These results demonstrate the superiority of our $\mathcal{A}LLM4ADD$ for the ADD task.
  • As the training data decreases from $ASV@full$ to $ASV@1/16$, the performance drop of $\mathcal{A}LLM4ADD$ is more moderate compared to that observed with end-to-end and conventional pipeline baselines. For example, $\mathcal{A}LLM4ADD^{\star}$ observes an EER increase from 1.30% to 3.13%, and $\mathcal{A}LLM4ADD^{\triangle}$ from 0.41% to 2.45%, whereas the AASIST model suffers a substantial increase in EER from 0.83% to 5.54%. We attribute this phenomenon to the ALLM’s ability to learn and apply transferable skills from limited data.
  • Training the audio encoder can often lead to improved performance. When averaged across $ASV@full$, $ASV@1/4$, $ASV@1/8$, and $ASV@1/16$ settings, the EER of $\mathcal{A}LLM4ADD^{\triangle}$ is 0.81% lower than that of $\mathcal{A}LLM4ADD^{\star}$. We speculate that this improvement stems from the encoder’s trainability, allowing it to extract more information relevant to audio authenticity.
@@ 6. Ablation Study
In this section, we further conduct extensive experiments to address the following research questions (Qs):
@@ 6.1. Generalization Capabilities of Models (Q1)
Table 4 illustrates that $\mathcal{A}LLM4ADD^{\triangle}$ outperforms both the end-to-end and conventional pipeline baselines on the In-the-Wild dataset. Specifically, when trained on $ASV@full$, $\mathcal{A}LLM4ADD^{\triangle}$ demonstrates an EER of 26.99%, whereas the best results from the end-to-end and pipeline models are 36.11% and 30.54%, respectively.
@@
$\mathcal{A}LLM4ADD^{\star}$ 32.04 65.28 73.82 45.12 40.66 51.64
$\mathcal{A}LLM4ADD^{\triangle}$ $\mathbf{26.99}$ $\mathbf{80.28}$ $\mathbf{90.83}$ $\mathbf{33.20}$ $\mathbf{51.84}$ 57.89

### v2

@@ Abstract.
Audio deepfake detection (ADD) has grown increasingly important due to the rise of high-fidelity audio generative models and their potential for misuse. Given that audio large language models (ALLMs) have made significant progress in various audio processing tasks, a heuristic question arises: Can ALLMs be leveraged to solve ADD?. In this paper, we first conduct a comprehensive zero-shot evaluation of ALLMs on ADD, revealing their ineffectiveness. To this end, we propose ALLM4ADD, an ALLM-driven framework for ADD. Specifically, we reformulate ADD task as an audio question answering problem, prompting the model with the question: “Is this audio fake or real?”. We then perform supervised fine-tuning to enable the ALLM to assess the authenticity of query audio. Extensive experiments are conducted to demonstrate that our ALLM-based method can achieve superior performance in fake audio detection, particularly in data-scarce scenarios. As a pioneering study, we anticipate that this work will inspire the research community to leverage ALLMs to develop more effective ADD systems. Code is available at https://github.com/ucas-hao/qwen_audio_for_add.git
@@ 1. Introduction
Recently, Audio Large Language Models (ALLMs) (Gong et al., 2024; Zhang et al., 2023b; Chu et al., 2023, 2024) have demonstrated remarkable progress across a wide range of audio processing tasks, including audio captioning (Tang et al., 2024; Shan et al., 2025; Wu et al., 2024c) and speech recognition (Li et al., 2024b; Fathullah et al., 2024; Chu et al., 2023). Notable models include Qwen-Audio (Chu et al., 2023), which integrates the Whisper encoder (Radford et al., 2023) with the text Qwen LLM (Bai et al., 2023), enabling the latter to understand audio. However, their performance on ADD task remains unexplored. This raises a critical question: Can ALLMs effectively address the ADD task?
To answer this question, we present pioneering work that leverages ALLMs for ADD task. To the best of our knowledge, this is the first paper to tackle ADD using the ALLM-based approach. First, we conduct a comprehensive quantitative evaluation of ALLM’s capabilities in ADD task. Our experimental results reveal that existing ALLMs perform poorly in zero-shot fake audio detection, primarily due to the mismatch between their pretraining objectives and the fake audio detection requirements. To enhance their performance for fake audio detection, we further propose a novel framework called ALLM4ADD. Specifically, we reformulate ADD task as an Audio Question Answering (AQA) problem, prompting the model with the question “Is this audio fake or real?” and instructing it to generate the correct answer. We then employ supervised fine-tuning (SFT) to endow the ALLMs the capability to answer “Fake” if the query audio is fake, and conversely, “Real” if it is real. Extensive experimental results demonstrate that ALLM4ADD achieves superior performance compared to existing conventional pipeline and end-to-end models, particularly in data-scarce scenarios. These findings collectively highlight the advantages of our ALLM-based approach for fake audio detection.
@@
  • We propose ALLM4ADD, a novel framework which reformulates ADD task as an AQA problem, to successfully endow ALLMs with fake audio detection capability.
@@ 2.2. Audio Large Language Model
In the past year, modern Large Language Models (LLMs) have demonstrated powerful reasoning and understanding abilities (Touvron et al., 2023; Bai et al., 2023; Wu et al., 2024a, b; Yu et al., 2025). To extend the application scope of LLMs beyond pure text tasks, many LLM-based multimodal models (Liu et al., 2023a; Bai et al., 2025; Zhang et al., 2023b; Gong et al., 2024; Zhang et al., 2023a; Sun et al., 2025) have been developed.
@@
While existing ALLMs have achieved notable success in tasks such as audio caption (Tang et al., 2024; Shan et al., 2025; Wu et al., 2024c) and speech recognition (Li et al., 2024b; Fathullah et al., 2024; Chu et al., 2023), their application in detecting fake audio remains unexplored. Therefore, in this paper, we formulate audio deepfake detection as an audio question answering task, leveraging the advanced understanding capabilities of these ALLMs to address this emerging challenge.
3. Can ALLMs detect fake audio zero-shot?
@@ 4.1. Task Formulation
In order to take advantage of the Audio Large Language Model (ALLM), ALLM4ADD formulates the audio deepfake detection (ADD) task as a audio question answering (AQA) problem. In this framework, the input comprises two crucial components: a query audio $A$ that needs to be classified as real or fake and an instruction prompt $q$, which guides ALLM4ADD in its analysis of the query audio. The instruction $q$ can take on various forms (e.g., “Is this audio fake or real?”). The output of this framework corresponds to the answer text $y$. While $y$ can be any text in principle, we constrain it to two options: ”Fake” and ”Real” during training, aligning with the ground truth to the original binary classification problem.
@@ 4.2. Model Architecture
As shown in Figure 1, ALLM4ADD adopts an architecture similar to the Qwen-Audio series, consisting of an audio encoder and a large language model. This framework takes two inputs: a query audio and a natural language instruction. Below, we describe each sub-module in detail.
@@
Our ALLM4ADD adopts a LLM as its foundation component. The model is initialized with pre-trained weights from Qwen-audio-chat default, which is a 32-layer Transformer decoder model with a hidden size of 4096, encompassing a total of 7.7B parameters. The LLM is employed to process audio representations and corresponding instructions, subsequently generating responses capable of discerning authenticity. The format input to the LLM follows this format:
@@ 4.3. Supervised Fine-tuning
Additionally, we categorize our methods based on the trainability of the audio encoder: $\mathbf{ALLM4ADD^{\star}}$ denotes that the encoder is frozen, while $\mathbf{ALLM4ADD^{\triangle}}$ indicates that the encoder is trainable. Our motivation is to transform the feature space of the audio embeddings into one that can effectively discriminate between real and fake by training the audio encoder. This strategy aims to capture more detailed nuances of audio authenticity, thereby enhancing the model’s performance in detecting fake audio.
@@ 5.1.2. Baselines
For end-to-end models, we select five competitive methods that provide open-source code: RawNet2 (Tak et al., 2021b), AASIST (Jung et al., 2021), RawGAT-ST (Tak et al., 2021a), Rawformer (Liu et al., 2023b) and RawBMamba (Chen et al., 2024b). RawNet2 (Tak et al., 2021b) is a convolutional neural network operating directly on raw audio waveforms, utilizing residual blocks and Sinc-Layers (Ravanelli and Bengio, 2018) as band-pass filters for effective ADD. AASIST (Jung et al., 2021) employs a heterogeneous stacking graph attention layer to model artifacts across temporal and spectral segments. RawGAT-ST (Tak et al., 2021a) utilizes spectral and temporal sub-graphs integrated with a graph pooling strategy, effectively processing complex auditory environments. Rawformer (Liu et al., 2023b) integrates convolution layer and transformer to model local and global artefacts and relationship directly on raw audio. RawBMamba (Chen et al., 2024b) proposes an end-to-end bidirectional state space model to capture both short- and long-range discriminative information.
@@ 5.1.3. Evaluation Metric
Table 3. Performance comparison of our methods with conventional pipeline and end-to-end models. (Train) and (Frozen) represent whether the self-supervised features are trainable. $\mathbf{ALLM4ADD^{\star}}$ and $\mathbf{ALLM4ADD^{\triangle}}$ denote the audio encoder is frozen and trainable, respectively. We train the models across: ASV@full, ASV@1/4, ASV@1/8, and ASV@1/16, and report EER (), ACC (), and AUC () on the ASVspoof2019 LA evaluation set. Best results in each column are highlighted in bold.
@@
$ALLM4ADD^{\star}$ 1.30 99.26 99.88 1.97 98.87 99.65 2.46 $\mathbf{98.69}$ 99.54 3.13 98.11 $\mathbf{99.43}$
$ALLM4ADD^{\triangle}$ $\mathbf{0.41}$ $\mathbf{99.39}$ $\mathbf{99.97}$ $\mathbf{1.15}$ $\mathbf{99.12}$ $\mathbf{99.71}$ $\mathbf{1.63}$ 98.55 $\mathbf{99.75}$ $\mathbf{2.45}$ $\mathbf{98.52}$ 99.39
@@ 5.2. Implementation Details
We use Qwen-audio-chat weights as our default initial weights. We optimize our model using the Adam optimizer with hyperparameters $\beta=(0.9,0.95)$ and implement a cosine learning rate scheduler. The warm-up ratio is set at 0.01 for $ASV@full$, $ASV@1/4$, and $ASV@1/8$ configurations, and 0.05 for $ASV@1/16$. We investigate two versions of our ALLM4ADD: $ALLM4ADD^{\star}$ and $ALLM4ADD^{\triangle}$. For $ALLM4ADD^{\star}$, the audio encoder is frozen; for $ALLM4ADD^{\triangle}$, the audio encoder is trainable. The initial learning rate is determined through beam search within the range of [3e-5, 4e-5, 5e-5, 1e-4]. Additionally, we apply a weight decay of 0.1 and a gradient clipping threshold of 1.0 to maintain training stability. Training epochs varies: 5 epochs for $ASV@full$, $ASV@1/4$, and $ASV@1/8$, and increase to 10 epochs for $ASV@1/16$. Furthermore, our training processes utilize bfloat16 precision with automatic mixed precision for efficiency. We conduct all experiments on a single Nvidia A100. For LoRA configuration, we configure our LoRA hyperparameter as follows: LoRA rank $r$ as 64, LoRA alpha (scaling factor) as 16 and LoRA dropout as 0.05. We apply LoRA to the query, key, value, and output projection layers.
For conventional pipeline baselines, we adhere to the hyperparameter provided in (Wang and Yamagishi, 2022). For end-to-end baselines, we adhere to the official codebase and train the models for 100 epochs. To maintain a fair comparison across all methods, we do not employ data augmentation techniques.
@@ 5.3. Experimental Results
To demonstrate the superiority of our ALLM-based fake audio detection method, we compare it with extensive baselines detailed in Sec. 5.1.2. The models are trained on $ASV@full$, $ASV@1/4$, $ASV@1/8$, and $ASV@1/16$, and are evaluated on the ASVspoof2019 evaluation set. From Table 3, we can draw the following observations:
  • Our ALLM4ADD can achieve excellent results in ADD task. When trained on $ASV@full$, $ALLM4ADD^{\triangle}$ achieves an EER of 0.41%, surpassing the best performances of end-to-end and conventional pipeline baselines, which are 0.83% and 1.94%, respectively. Furthermore, ALLM4ADD maintains strong performance when data is scarce. Specifically, under the $ASV@1/16$ setting, $ALLM4ADD^{\triangle}$ still achieves an EER of 2.45% and an accuracy of 98.52%. These results demonstrate the superiority of our ALLM4ADD for the ADD task.
  • As the training data decreases from $ASV@full$ to $ASV@1/16$, the performance drop of ALLM4ADD is more moderate compared to that observed with end-to-end and conventional pipeline baselines. For example, $ALLM4ADD^{\star}$ observes an EER increase from 1.30% to 3.13%, and $ALLM4ADD^{\triangle}$ from 0.41% to 2.45%, whereas the AASIST model suffers a substantial increase in EER from 0.83% to 5.54%. We attribute this phenomenon to the ALLM’s ability to learn and apply transferable skills from limited data.
  • Training the audio encoder can often lead to improved performance. When averaged across $ASV@full$, $ASV@1/4$, $ASV@1/8$, and $ASV@1/16$ settings, the EER of $ALLM4ADD^{\triangle}$ is 0.81% lower than that of $ALLM4ADD^{\star}$. We speculate that this improvement stems from the encoder’s trainability, allowing it to extract more information relevant to audio authenticity.
@@ 6. Ablation Study
This section investigates the following research questions (Qs).
@@ 6.1. Generalization Capabilities of Models (Q1)
Table 4 illustrates that $ALLM4ADD^{\triangle}$ outperforms both the end-to-end and conventional pipeline baselines on the In-the-Wild dataset. Specifically, when trained on $ASV@full$, $ALLM4ADD^{\triangle}$ demonstrates an EER of 26.99%, whereas the best results from the end-to-end and pipeline models are 36.11% and 30.54%, respectively.
@@
$ALLM4ADD^{\star}$ 32.04 65.28 73.82 45.12 40.66 51.64
$ALLM4ADD^{\triangle}$ $\mathbf{26.99}$ $\mathbf{80.28}$ $\mathbf{90.83}$ $\mathbf{33.20}$ $\mathbf{51.84}$ 57.89